# From Continual Learning to SGD and Back: Better Rates for Continual Linear Models

### v2

@@ Appendix B Auxiliary Proofs
In a realizable setting (Assumption 2.2), the iterates of Scheme 1 under a random task ordering $\tau$ (with or without replacement) hold $\forall k\geq 1$,
@@
  • Proof.
    As discussed in Section 3.1, Scheme 2 governs the updates of the iterates $\mathbf{w}_{t}\in\mathbb{R}^{d}$. Under Assumption 2.2, we define the orthogonal projection as ${\mathbf{P}}_{\tau(t)}\triangleq\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}$, revealing a recursive form:
@@
    [Assumption 2.2]delimited-[]Assumption $\end{subarray}\right]$ $=\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\mathbf{w}_{\star}+\left(\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\right)\mathbf{w}_{t-1}=(\mathbf{I}_{d}-{\mathbf{P}}_{\tau\left({t}\right)})\mathbf{w}_{\star}+{\mathbf{P}}_{\tau\left({t}\right)}\mathbf{w}_{t-1}$
@@
    $\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq\frac{1}{k}\sum_{t=1}^{k}\bigg{(}~{}\mathbb{E}_{\tau}\underbrace{\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}}_{\leq R^{2}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\text{, since projections contract}}+$
@@
    $\leq\frac{1}{k}\sum_{t=1}^{k}\bigg{(}~{}R^{2}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}+$
@@
    For the POCS case, simply replace $\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$ with $\left\|\mathbf{w}_{k}-\bm{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2}$. $\blacksquare$
@@
      Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}$
@@
      Gradient:       $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$
@@
      Finally, the convexity of $f_{m}$ is immediate since $\nabla_{\mathbf{w}}^{2}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\succeq\mathbf{0}$. $\blacksquare$
@@
    Here, we prove the main result, followed by the necessary auxiliary corollaries and lemmas in Section C.1.
@@
    This map plays a central role in our analysis and has been studied in similar forms in prior work [23]. Note that ${\mathbf{P}}_{m}$ is an orthogonal projection, which implies that it is symmetric and idempotent. Thus,
@@
    where $Q^{t}$ denotes $t$ applications of $Q$. The map $Q$ captures the evolution of the error’s second-moment under Kaczmarz updates, offering a tractable approach to analyzing the algorithm’s convergence. The expected loss at step $t$ is given by
@@
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\left\|{\mathbf{X}\mathbf{z}_{t}}\right\|^{2}}\right]$
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\mathbf{z}_{t}^{\top}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right)}\right]$
    $=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]}\right)=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]}\right).$
@@ C.1 Key Properties and Auxiliary Lemmas
  • Proof.
    Let $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, for all $i\in\left[{T}\right]$, $\mathbf{0}\preccurlyeq{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}$. Meaning $Q\left[{\mathbf{A}}\right]$ is PSD as a convex combination of PSD matrices. $\blacksquare$
@@
  • Proof.
    For $n=1$, given by Corollaries C.3 and C.4. For $n>1$, this follows trivially by induction. $\blacksquare$
@@
  • Proof.
    Since $\mathbf{0}\preccurlyeq\mathbf{C}=\mathbf{C}^{\top}$, it has a square symmetric PSD root $\mathbf{C}^{1/2}$. Given that $\mathbf{A},\mathbf{B}$ are symmetric and $\mathbf{A}\preccurlyeq\mathbf{B}$, it follows that $\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}\preccurlyeq\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}$ [from 28, Theorem 7.7.2.a]. Applying the cyclic property of the trace and using the fact that for symmetric matrices ordered in the Löwner sense, their traces are also ordered [28, Corollary 7.7.4.d], we obtain
@@
    [Corollary C.4]delimited-[]Corollary $\end{subarray}\right]$ $=\operatorname{tr}\left({Q\left[{\mathbf{B}}\right]^{\top}\mathbf{A}}\right)=\operatorname{tr}\left({\mathbf{A}^{\top}Q\left[{\mathbf{B}}\right]}\right)=\langle\mathbf{A},Q\left[{\mathbf{B}}\right]\rangle\,.$
    $\blacksquare$
@@
    From the Rayleigh quotient characterization of eigenvalues, this implies that every eigenvalue $\lambda$ of $Q$ satisfies $0\leq\lambda\leq 1$, i.e., $\sigma(Q)\subset[0,1]\,$. $\blacksquare$
@@
    Therefore, $\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en}$. $\blacksquare$
@@
$\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}\leq\min\left({\sqrt{T\bar{r}},\sqrt{d-\bar{r}}}\right)$
  • Proof.
@@
[…]
    and therefore $\mathbb{E}\mathcal{L}(\mathbf{w}_{k})\leq\frac{2R^{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}$, which proves the first claim. The second claim follows immediately from Lemma B.1, and we are done. $\blacksquare$
@@
For the second, parameter independent term, note that from Lemma 3.1, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$, and recall that the iterates $\mathbf{w}_{t}$ follow the SGD dynamics with $\eta=1$ (Reduction 2). Hence, by Lemma E.5, without-replacement SGD with $\beta=\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@ Appendix E Proofs of Last-Iterate SGD Bounds (Section 5)
In this section we provide proofs and full technical details of our upper bounds for least squares SGD. We begin by recording a few elementary well-known facts, which can be found in e.g., Bubeck [8]. We provide proof for completeness.
@@
    $\iff\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})$ $=\frac{1}{2\eta}\left(\left\|{\mathbf{w}_{t}-\tilde{\mathbf{w}}}\right\|^{2}-\left\|{\mathbf{w}_{t+1}-\tilde{\mathbf{w}}}\right\|^{2}\right)+\frac{\eta}{2}\left\|{\mathbf{g}_{t}}\right\|^{2}.$
@@
    $\sum_{t=0}^{T}\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})$ $=\frac{1}{2\eta}\left(\left\|{\mathbf{w}_{0}-\tilde{\mathbf{w}}}\right\|^{2}-\left\|{\mathbf{w}_{T+1}-\tilde{\mathbf{w}}}\right\|^{2}\right)+\frac{\eta}{2}\sum_{t=0}^{T}\left\|{\mathbf{g}_{t}}\right\|^{2}$
@@
    which completes the proof. $\blacksquare$
@@
    The second claim now follows by using the fact that $f(\mathbf{w}_{\star})\leq f(\mathbf{w}^{+})$. $\blacksquare$
@@ E.1 Proofs for With Replacement Orderings
    and the result follows after dividing by $(2-\eta\beta)$. $\blacksquare$
@@
    [Lemma E.3]delimited-[]Lemma $\end{subarray}\right]$ $\leq\sum_{t=T-k}^{T}\left\langle\mathbf{g}_{t},\mathbf{w}_{t}-\mathbf{z}\right\rangle+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
@@
    $\mathbb{E}f_{T}(\mathbf{w}_{T})$ $=\left({eT}\right)^{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\cdot\frac{1}{T+1}\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\,.$
    $\blacksquare$
@@ E.2 Extending the SGD Bounds to Without Replacement Orderings
The proof, given next, is based on the algorithmic stability of SGD [6, 61, 25], and more specifically, on a variant of stability, suitable for without replacement sampling [63, 35].
The proof of our theorem follows by a combination of two lemmas. The first, stated below, establishes a bound on the expected “next sample” loss and follows immediately by combining Lemmas 5.3 and 5.2 (notice that $\eta<\frac{2}{\beta}\Longrightarrow\exp\big{(}{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\big{)}\mapsto\exp\big{(}{z\big{(}{1-\frac{z}{4}}\big{)}}\big{)}$ for $z\in(0,2)$, which is monotonic increasing and upper bounded by $e$).
@@
For any step-size $\eta<2/\beta$ and initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, without-replacement SGD Eq. 7 satisfies, for all $11\leq T\leq n-$ 1:
@@
The proof of Lemma E.6 builds on an algorithmic stability argument similar to that given in Lei and Ying [38], combined with the without-replacement stability framework proposed by Sherman et al. [63]. Before turning to the proof given in the next subsection, we quickly prove Theorem E.4.
  • Proof of Theorem E.4.
    By Lemmas E.6 and E.5,
@@
$\mathbf{w}_{\tau}^{\pi}$ $\triangleq\text{The iterate of SGD on step }\tau\text{ when run on permutation }\pi.$
@@
$\mathbf{w}_{\tau}^{\pi(i\leftrightarrow t)}$ $\triangleq\text{The iterate of SGD on step }\tau\text{ when run on }\pi(i\leftrightarrow t).$
@@
    the result follows. $\blacksquare$
Our next lemma, originally given in Sherman et al. [63, Lemma 2 therein], can be thought of as a without-replacement version of the well known stability $\iff$ generalization argument of the with-replacement sampling case [61, 25].
@@
    where the last equality is immediate since by definition, $\pi(i\leftrightarrow t)_{i}=\pi_{t}$. The claim now follows by linearity of expectation. $\blacksquare$
@@
    Next, we bound $\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$. For any $10\leq\tau\leq t-$ 1, we denote $f_{\tau}\triangleq f(\cdot;\pi_{\tau})$, and $f_{\tau}^{(i\leftrightarrow t)}\triangleq f(\cdot;\pi(i\leftrightarrow t)_{\tau})$. Observe that for any $\tau$ such that $\tau\neq i$, we have $f_{\tau}=f_{\tau}^{(i\leftrightarrow t)}$, thus, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in 25):
@@
    which, when combined with Proof of Lemma E.6. yields:
@@
    which completes the proof. $\blacksquare$

### v3

@@ Appendix B Auxiliary Proofs
In a realizable setting (Section 2), the iterates of Scheme 1 under a random task ordering $\tau$ (with or without replacement) hold $\forall k\geq 1$,
@@
Proof B.2.
As discussed in Section 3.1, Scheme 2 governs the updates of the iterates $\mathbf{w}_{t}\in\mathbb{R}^{d}$. Under Section 2, we define the orthogonal projection as ${\mathbf{P}}_{\tau(t)}\triangleq\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}$, revealing a recursive form:
@@
[…]
@@
$\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq\frac{1}{k}\sum_{t=1}^{k}\bigg(~\mathbb{E}_{\tau}\underbrace{\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}}_{\leq R^{2}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\text{, since projections contract}}+\$
@@
$\leq\frac{1}{k}\sum_{t=1}^{k}\bigg(~R^{2}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}+\$
@@
For the POCS case, simply replace $\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$ with $\left\|\mathbf{w}_{k}-\bm{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2}$.
@@
    Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}$ .
@@
    Gradient:    $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$ .
@@
Finally, the convexity of $f_{m}$ is immediate since $\nabla_{\mathbf{w}}^{2}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\succeq\mathbf{0}$.
@@
Here, we prove the main result, followed by auxiliary corollaries and lemmas in Section C.1.
@@
This map plays a central role in our analysis and has been studied in similar forms in prior work (Guo et al., 2022). Note that ${\mathbf{P}}_{m}$ is an orthogonal projection, i.e., symmetric and idempotent. Thus,
@@
where $Q^{t}$ denotes $t$ applications of $Q$. The map $Q$ captures the evolution of the error’s second-moment under Kaczmarz updates, offering a tractable approach to analyzing the algorithm’s convergence.
The expected loss at step $t$ is given by
@@
$=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\left\|{\mathbf{X}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\mathbf{z}_{t}^{\top}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}}\right]$
$=\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right)}\right]=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]}\right)=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]}\right).$
@@ C.1 Key Properties and Auxiliary Lemmas
Proof C.5.
Let $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, for all $i\in\left[{T}\right]$, $\mathbf{0}\preccurlyeq{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}$. Meaning $Q\left[{\mathbf{A}}\right]$ is PSD as a convex combination of PSD matrices.
@@
Proof C.9.
For $n=1$, given by Corollaries C.4 and C.6. For $n>1$, this follows trivially by induction.
@@
Proof C.11.
Since $\mathbf{0}\preccurlyeq\mathbf{C}=\mathbf{C}^{\top}$, it has a square symmetric PSD root $\mathbf{C}^{1/2}$. Given that $\mathbf{A},\mathbf{B}$ are symmetric and $\mathbf{A}\preccurlyeq\mathbf{B}$, it follows that $\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}\preccurlyeq\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}$ (from Horn and Johnson, 2012, Theorem 7.7.2.a). Applying the cyclic property of the trace and using the fact that for symmetric matrices ordered in the Löwner sense, their traces are also ordered (Horn and Johnson, 2012, Corollary 7.7.4.d), we obtain
@@
[…]
@@
From the Rayleigh quotient characterization of eigenvalues, this implies that every eigenvalue $\lambda$ of $Q$ satisfies $0\leq\lambda\leq 1$, i.e., $\sigma(Q)\subset[0,1]\,$.
@@
Therefore, $\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en}$.
@@
$\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}\leq\min\left({\sqrt{T\bar{r}},\sqrt{d-\bar{r}}}\right)$ .
Proof C.23.
@@
$\mathbb{E}\bar{f}(\mathbf{w}_{k})\leq\frac{e\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\sqrt[4]{k}}\leq\frac{2\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}\,$,
and therefore $\mathbb{E}\mathcal{L}(\mathbf{w}_{k})\leq\frac{2R^{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}$, which proves the first claim. The second claim follows immediately from Lemma B.1, and we are done.
@@
For the second, parameter independent term, note that from Lemma 3.1, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$, and recall that the iterates $\mathbf{w}_{t}$ follow SGD dynamics with $\eta=1$ (Reduction 2). Hence, by Lemma E.10, without-replacement SGD with $\beta=\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@ Appendix E Proofs of Last-Iterate SGD Bounds (Section 5)
In this section we provide proofs and full technical details of our upper bounds for least squares SGD. We begin by recording a few elementary well-known facts, which can be found in e.g., Bubeck (2015). We provide proof for completeness.
@@
[…]
@@
[…]
@@
which completes the proof.
@@
The second claim now follows by using the fact that $f(\mathbf{w}_{\star})\leq f(\mathbf{w}^{+})$.
@@ E.1 Proofs for With Replacement Orderings
and the result follows after dividing by $(2-\eta\beta)$.
@@
[…]
@@
$\mathbb{E}f_{T}(\mathbf{w}_{T})$ $\leq\left({eT}\right)^{\eta\beta\big({1-\frac{\eta\beta}{4}}\big)}\cdot\frac{1}{T+1}\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\,.$
@@ E.2 Extending the SGD Bounds to Without Replacement Orderings
The proof, given next, is based on the algorithmic stability of SGD (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010; Hardt et al., 2016), and more specifically, on a variant of stability, suitable for without replacement sampling (Sherman et al., 2021; Koren et al., 2022).
The proof of our theorem follows by a combination of Lemmas E.10 and E.11. The first, stated below, establishes a bound on the expected “next sample” loss and follows immediately by combining Lemmas 5.3 and 5.2 (notice that $\eta<\frac{2}{\beta}\Longrightarrow\exp\big({\eta\beta\big({1-\frac{\eta\beta}{4}}\big)}\big)\mapsto\exp\big({z\big({1-\frac{z}{4}}\big)}\big)$ for $z\in(0,2)$, which is monotonic increasing and upper bounded by $e$).
@@
For any step-size $\eta<2/\beta$ and initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, without-replacement SGD Eq. 7 satisfies, for all $1\leq T\leq n-1$:
@@
The proof of Lemma E.11 builds on an algorithmic stability argument similar to that given in Lei and Ying (2020), combined with the without-replacement stability framework proposed by Sherman et al. (2021). Before turning to the proof given in the next subsection, we quickly prove Theorem E.9.
Proof E.12 (of Theorem E.9).
By Lemmas E.11 and E.10,
@@
$\pi(j\leftrightarrow k)$ $\triangleq\pi\text{ after swapping the }j\textsuperscript{th}\text{ and }k\textsuperscript{th}\text{ coordinates},$
@@
[…]
@@
the result follows.
Our next lemma, originally given in Sherman et al. (2021, Lemma 2 therein), can be thought of as a without-replacement version of the well known stability $\iff$ generalization argument of the with-replacement sampling case (Shalev-Shwartz et al., 2010; Hardt et al., 2016).
@@
where the last equality is immediate since by definition, $\pi(i\leftrightarrow t)_{i}=\pi_{t}$. The claim now follows by linearity of expectation.
@@
Next, we bound $\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$. For any $0\leq\tau\leq t-1$, we denote $f_{\tau}\triangleq f(\cdot;\pi_{\tau})$, and $f_{\tau}^{(i\leftrightarrow t)}\triangleq f(\cdot;\pi(i\leftrightarrow t)_{\tau})$. Observe that for any $\tau$ such that $\tau\neq i$, we have $f_{\tau}=f_{\tau}^{(i\leftrightarrow t)}$, thus, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
@@
which, when combined with E.17 yields:
@@
which completes the proof.